OpenStudy (anonymous):

What is -7/9 times 3/5

OpenStudy (anonymous):

Help medal to whoever helps

OpenStudy (jango_in_dtown):

(-7/9).(3/5)=-7/(3.5)=-7/15

OpenStudy (anonymous):

Thanks

OpenStudy (jango_in_dtown):

@Jersihelp16